# Dialectal and Low Resource Machine Translation for Aromanian

### v1

@@ 4 Aromanian Sentence Embeddings
Throughout the data collection process, we used a sentence embedder compatible with Aromanian to align documents. Thus, we fine-tune a language-agnostic Bert Sentence Encoder (LaBSE) Feng et al. (2020) model with an extended vocabulary, using a similar approach to Dale (2022). The purpose is to embed Aromanian sentences in the same latent space as the sentences from Romanian and English.
First, we extend the tokenizer’s vocabulary on a monolingual Aromanian corpus (i.e., we utilize Aromanian sentences from an intermediate version of the corpus). We perform this extension with the BPE algorithm Sennrich (2015), and we obtain k new tokens. This lowers the tokens per word ratio of LaBSE tokenizers from to .
Secondly, we fine-tune the model on both Aromanian-Romanian and Aromanian-English parallel pairs, updating only the embedding matrix, with the contrastive loss, in a similar manner to Feng et al. (2020).
At each training step, we randomly choose a batch of pairs in a random pair of languages out of the two established language pairs. We calculate the dot product of every possible sentence pair’s embeddings and only reward the pairs that are supposed to match. However, we penalize the matching pairs by substracting a small margin , to avoid over-fitting. We stop after 150k steps, as the graphs of both loss and accuracy (of matching translated sentences) seem to flatten after this point (Figure 2). Thus, in a batch of random sentence pairs, the model pairs each Aromanian sentence with its corresponding translation with an accuracy of more than . The training hyper-parameters are available in Appendix C.
The model can be used in multiple ways: (1) to calculate the BERTScore Zhang et al. (2019) and validate our MT models (see subsection 6.1); and (2) to align sentences in parallel documents. Similarly to Dale (2022), we use dynamic programming to select a sequence of sentence pairs that have the highest possible sum of sentence similarity scores. This method has been used to mine parallel sentences in section 3 from various genres.
@@ 5.1 GPT Baseline
Inspired by recent MT results Hendy et al. (2023); Kocmi et al. (2024) where closed-source systems achieved state-of-the art performance on machine translation, we employ the GPT-4o model, in hopes of leveraging its extensive multilingual pre-training Achiam et al. (2023). Additionally, we suspect that the model had been exposed to Aromanian in its training stage. For example, the Aromanian Wikipedia101010https://roa-rup.wikipedia.org, Accessed: 2024-09-16 although of low quality, is part of the Common Crawl111111https://commoncrawl.org, Accessed: 2024-09-16. We use OpenAI’s gpt-4o model in zero-shot mode to translate all pairs in our test split, which we later evaluate in section 6.
@@
| model | ron-¿rup | rup-¿ron | eng-¿rup | rup-¿eng | eng-¿ron | ron-¿eng |
@@
| Qwen2 7B Instruct | 42.1 | 48.9 | 35.1 | 50.9 | 56.6 | 75.7 |
| RoLLaMA | 42.1 | 48.9 | 35.9 | 51.3 | 58.4 | 76.9 |
| TowerInstruct v0.2 | 41.9 | 49.1 | 27.9 | 43.2 | 48.6 | 74.8 |
| LLaMA 3.1 8B Instruct | 42.4 | 49.6 | 35.2 | 51.7 | 58.9 | 76.9 |
| NLLB 1.3B | 43.7 | 51.5 | 39.5 | 53.9 | 61.5 | 75.9 |
| NLLB 600M | 45.0 | 51.0 | 40.4 | 53.0 | 60.1 | 75.1 |
@@ 5.2 NLLB Machine Translation
Seeking to benefit from transfer learning, we aim to fine-tune NLLB-200 Costa-jussà et al. (2022), an encoder-decoder transformer architecture model that can translate between any of its 202 languages. Our final goal is a model that has the ability to translate between Aromanian, Romanian, and English. The last two are already supported by the NLLB.
The NLLB tokenizer uses language tags, i.e. special tokens are added to the source and target texts, used by the model in the pre-training phase for identifying the source and target languages. The special tokens for Romanian and English are <ron_Latn> and <eng_Latn> respectively. We expand the tokenizer vocabulary with the <rup_Latn> token for Aromanian. Its meaning is supposedly constructed during the fine-tuning step. To achieve this, we initialize the embedding value of the newly added token as the mean of the <ron_Latn> and <ell_Grek> (Greek) embedding vectors. This is justified by the close relationship of Aromanian with standard Romanian and the large proportion of Greek influence and loanwords Dragomirescu (2020).
We experiment with the 1.3B and 600M distilled models of the original 3.3 billion parameter model. We train for k steps, where each sentence pair in the batch is in a random direction between Aromanian, Romanian, and English (i.e. six possible directions), aiming to better stabilize the training process this way. Also note that NLLB can already translate between Romanian and English. The hyper-parameters used for training the two models are available in Appendix D.
For later testing (section 6), we select the NLLB-600M checkpoint at k training steps, and the NLLB-1.3B checkpoint at k training, as these produce the highest BLEU scores on the dev split.
@@ 5.3 LLMs for Machine Translation
LLMs have proven useful in numerous NLP tasks, including machine translation. Ranging from zero-shot or few-shot translation Kocmi et al. (2024), to fine-tuning LLMs for MT Alves et al. (2024).
We choose to perform full fine-tuning on four instruction-tuned LLMs (i.e., base LLMs fine-tuned for instruction following). The models are as follows:
LLaMA 3.1 8B Instruct Dubey et al. (2024) - the most recent multilingual LLaMA model covering a diverse set of instructions; Romanian appears to be incidentally covered in the 15% of multilingual tokens given during pretraining
Qwen2 7B Instruct Yang et al. (2024) - a multilingual model from Alibaba Cloud covering as many as 30 languages, including 4 well resourced Western and Italo-Romance languages: French, Spanish, Portuguese, Italian
RoLLaMA 3 8B Instruct Masala et al. (2024) - a LLaMA3-based model fine-tuned specifically to respond to tasks and instructions in Romanian; the exposure of this model to Romanian should allow for an easier transfer to Aromanian, its Eastern Romance sibling
TowerInstruct 7B v0.2 Alves et al. (2024) - a LLaMA2-based model trained specifically to solve translation tasks (document-level translation, terminology-level translation, etc.) on 10 languages, including 4 well resourced Western and Italo-Romance languages: French, Spanish, Portuguese, Italian
@@
As with NLLB, we have the same objective of translating between Aromanian, Romanian, and English, so we train the LLMs with samples from all six possible directions. In the case of the four models, fine-tuning for more than epoch led to an increased loss on the dev split. Thus, we keep only the checkpoints fine-tuned for epoch with sample packing Krell et al. (2021). An example of a prompt for the LLaMA 3.1 model can be found in Table 3. More details about each model’s training hyper-parameters and prompt format can be found in Appendix E.
@@ 6.1 Automatic evaluation
| Participant | Direction | Fluency | Style | Meaning | |||
| HT | MT | HT | MT | HT | MT | ||
| Subject 1 | ron->rup | 9.66 | 9.73 | 9.60 | 9.60 | 9.73 | 9.86 |
| rup->ron | 9.46 | 9.66 | 9.46 | 9.53 | 9.46 | 9.33 | |
| Subject 2 | ron->rup | 9.66 | 9.66 | 9.33 | 9.60 | 9.66 | 9.73 |
| rup->ron | 8.13 | 9.0 | 7.73 | 8.73 | 8.2 | 8.93 | |
We evaluated all our models in all possible directions between Aromanian, Romanian, and English. We note here that the English references are machine translated and the results involving English should be taken with reservation. However, we provide the evaluation scores here for completeness.
The automatic metrics that we employ are BLEU Papineni et al. (2002), ChrF++ Popović (2015), and BERTScore Zhang et al. (2019). We report the ChrF++ scores for all the models in Table 4. All other results are available in Appendix F.
Overall, for English-Romanian pairs, GPT-4o appears to perform better than our models trained for the same language pairs. This could be an indicator that GPT-4o generates similar results to Google Translate, however, we cannot derive any conclusions at this point with respect to quality.
More importantly, the scores for the Aromanian-Romanian language pairs obtained by our models are considerably higher than those obtained by the GPT-4o model. Among our entire model set, the best performing are the NLLB family, obtaining consistently higher scores than their fine-tuned instruction LLM counterparts, regardless of the automatic metric (Appendix F). When translating into Aromanian (-¿rup), the differences are considerable - in every single case NLLB-type models perform significantly better than LLM-based translation models. However, when translating into a well-resourced language such as English or Romanian, the differences between NLLB models and LLM are less significant, only a few points apart.
All types of models struggle more to generate good quality Aromanian, a fact visible also in the overall lower BLEU scores (max 17 ron-¿rup, 13.4 eng-¿rup). A manual inspection of the translations reveals that the models tend to translate sentences word-by-word which misses word overlaps in BLEU-based metrics for out-of-domain evaluation on mixed-genre literary, news, and essay. More details on manual evaluation are presented in the next section.
@@ 6.2 Human Evaluation
Two human evaluators were involved in this study. Both were native Aromanian speakers residing in Romania, so they can speak both Romanian and Aromanian. We do not evaluate for English here. The sentences given for evaluation contain both human and machine translations in equal proportion and so each annotator evaluates a total of samples. For machine translation, we use our NLLB 600M fine-tuned model.
Annotators are instructed to assign direct assessment scores Graham et al. (2013) from 1 to 10 to evaluate the quality of each translation (1 being the lowest score, 10 being the highest). This scale is culturally motivated by the fact that it is familiar to the study participants because it is used in Romanian schools at all levels of post-primary school.
They assigned a score taking into account three categories: fluency, style, and meaning (logical sense). Fluency refers to how grammatically correct the sentence is. Style denotes how likely it is for the speaker to build the sentence they are evaluating. Meaning alludes to the logical sense of the sentence.
The evaluation was carried out bilingually for all language pairs. The results can be seen in Table 5. The outcome is dubious, as it may appear that machine translation is rated better than human translation. Manually analyzing the samples, we observe that oftentimes, the human translations are more creative, whilst the model tends to keep the structure of the original sentence mot-a-mot. In addition, the small number of human evaluators together with the small sample size renders the human evaluation inconclusive.
@@ 7 Online translation system
Seeking to make our machine translation model accessible to the public and raise awareness of the endangered status of the Aromanian language, we deploy a translation system online. Our GUI is lightweight, easy to use, and multilingual, with implemented accessibility features for copying text content and switching between source and target language (Figure 3).
| ron->rup | rup->ron | |||
| Model | BLEU | ChrF++ | BLEU | ChrF++ |
| Unquantized | 17.3 | 45.03 | 30.95 | 51.01 |
| Quantized | 17.01 | 44.84 | 30.21 | 50.16 |
@@ 7.1 Quantization
We choose to deploy on bare-metal CPU the 600M distilled NLLB version of our translator, as it scores best on a range of categories and has the smallest number of parameters. We quantize it using the ctranslate2 Klein et al. (2020) engine to bit-integer weights. Additionally, the engine also applies other optimizations such as layer fusion, batch reordering, etc.
Regarding the performance loss that is expected with quantization, we report minimal differences in the automatic metrics we deploy. These effects can be seen in Table 6 where we analyze the Aromanian-Romanian direction of our test split. Differences of less than BLEU point are measured.
Because all our models were fine-tuned for sentence-level translations, we split each user’s input into sentences and process it in parallel. Our system can translate approximately tokens per second.
7.2 Orthography Converter
@@ 8 Conclusions
Building a dialectal machine translation system for a low-resource language comes with several challenges - from the lack of standardization and small online presence of the language, to the diverse varieties specific to different countries. Given Aromanian’s close resemblance to Romanian, we show that it is possible to use pre-existing knowledge of Romanian both from LLM-like transformers and sequence-to-sequence models to build a MT system for Aromanian, Romanian, and English. Based on automatic evaluation, it appears that smaller Romanian-pre-trained seq-to-seq models are better translators than LLM-based English-centric or multilingual fine-tuned models (that have not been pre-trained on Romanian). Regarding human evaluation, due to the relatively small size of the community and lack of access to a large pool of native speakers, we had difficulties carrying out an extensive human evaluation. Future studies are mandatory to quantify the limitations of our MT system across different dialectal regions. With our multigenre dataset, we hope to bring Aromanian into the spotlight of MT and corpus-based linguistic research and contribute to preservation efforts for this endangered language.

### v2

@@ 4 Aromanian Sentence Embeddings
To embed Aromanian sentences in the same latent space as Romanian and English sentences a language-agnostic BERT Sentence Encoder Feng et al. (2022) is fine-tuned following the methodology described by Dale (2022).
First, the tokenizer’s vocabulary is trained on a monolingual Aromanian corpus, using all Aromanian sentences prior to alignment. This is performed with the same WordPiece tokenizer from BERT Devlin et al. (2019), yielding 4,400 new tokens. As a result, the tokens-per-word ratio of the LaBSE tokenizer decreases from to .
Secondly, the model is fine-tuned on both Aromanian-Romanian and Aromanian-synthetic English parallel pairs, updating only the pooled output corresponding to Aromanian embeddings using contrastive loss.
At each training step, a batch of sentence pairs is randomly selected from one of the two pairs of languages. The dot product of the embeddings for all possible sentence pairs in the batch is computed, rewarding only the pairs that are correct matches. To prevent overfitting, a small margin of is subtracted from the reward for matching pairs. Training stops after k steps, as both the loss and accuracy graphs (for matching translated sentences) flatten beyond this point (Figure 2). In a batch of random sentence pairs, the model pairs each Aromanian sentence with its corresponding translation with an accuracy of over . Details on the training hyper-parameters are provided in Appendix C.
The model is used in two ways: (1) to calculate the BERTScore Zhang et al. (2020) and evaluate trained machine translation models (see Section 6.1); and (2) to mine and align sentences in parallel documents. Similarly to Dale (2022), dynamic programming is used to select a sequence of sentence pairs that have the highest possible sum of similarity scores.
@@ 5.1 GPT Baseline
Inspired by recent MT results Hendy et al. (2023); Kocmi et al. (2024) where closed-source systems achieved state-of-the-art performance on machine translation, we employ the GPT-4o model, in hopes of leveraging its extensive multilingual pre-training Achiam et al. (2024). Additionally, it is likely that the model has been exposed to Aromanian in its pre-training stage. We use OpenAI’s gpt-4o model in zero-shot mode to translate all pairs in the test split, which we later evaluate in Section 6.
@@
| model | ronrup | rupron | engrup | rupeng | engron | roneng |
@@
| Qwen2 7B Instruct | 45.8 | 50.5 | 38.8 | 51.9 | 58.5 | 76.2 |
| TowerInstruct v0.2 | 45.7 | 50.9 | 31.2 | 44.6 | 50.8 | 75.2 |
| LLaMA 3.1 8B Instruct | 46.3 | 51.4 | 38.8 | 52.6 | 60.8 | 77.2 |
| RoLLaMA | 46.0 | 50.6 | 39.7 | 52.2 | 60.2 | 77.3 |
| NLLB 1.3B | 47.8 | 53.2 | 43.6 | 54.7 | 63.2 | 76.3 |
| NLLB 600M | 49.1 | 52.7 | 44.5 | 53.8 | 62.2 | 75.5 |
@@ 5.2 NLLB Machine Translation
Seeking to benefit from transfer learning Khiu et al. (2024), we fine-tune NLLB-200 Costa-jussà et al. (2022), an encoder-decoder transformer architecture model that can translate between any of its 202 languages. The goal is to obtain a model that has the ability to translate between Aromanian, Romanian, and English. The last two are already supported by the NLLB.
The NLLB tokenizer uses language tags, i.e. special tokens are added to the source and target texts, which are employed by the model in the pre-training phase to identify the source and target languages. The special tokens for Romanian and English are <ron_Latn> and <eng_Latn>, respectively. We expand the tokenizer vocabulary with the <rup_Latn> token for Aromanian. Its meaning is supposedly constructed during the fine-tuning step. The embedding value of the newly added token is set to the mean of the <ron_Latn> and <ell_Grek> (Greek) embedding vectors. This is justified by the close relationship of Aromanian with standard Romanian and the large proportion of Greek influence and loanwords Pascaru and Kahl (2017); Dragomirescu (2020).
We experiment with the 1.3B and 600M distilled models of the original Mixture-of-Experts 54B billion parameter model. Each model is trained for k steps, where each sentence pair in the batch is in a random direction between Aromanian, Romanian, and English (i.e. six possible directions). At every k steps, we save a checkpoint and evaluate the dev set. The NLLB-600M checkpoint at k training steps, and the NLLB-1.3B checkpoint at k training steps produce the highest evaluation scores on the dev split. The hyper-parameters used for training the two models are available in Appendix D.
@@ 5.3 LLMs for Machine Translation
LLMs have proven to be the state of the art for general machine translation, from zero-shot or few-shot translation Kocmi et al. (2024) to different fine-tuning strategiesAlves et al. (2024).
We perform full fine-tuning on four trained for instruction following:
LLaMA 3.1 8B Instruct Grattafiori et al. (2024) – a multilingual model from Meta AI in which Romanian is incidentally covered in the of multilingual tokens used during pre-training
Qwen2 7B Instruct Yang et al. (2024) - a multilingual model from Alibaba Cloud covering as many as 30 languages, including 4 well-resourced Western and Italo-Romance languages: French, Spanish, Portuguese, Italian
RoLLaMA 3 8B Instruct Masala et al. (2024) – a LLaMA3-based model fine-tuned specifically to respond to tasks and instructions in Romanian, hoping to facilitate knowledge transfer to Aromanian
TowerInstruct 7B v0.2 Alves et al. (2024) – a LLaMA2-based model trained specifically to solve translation tasks (document-level translation, terminology-level translation, etc.) on 10 languages, including 4 well-resourced Western and Italo-Romance languages: French, Spanish, Portuguese, Italian
| Participant | Direction | Fluency | Style | Meaning | |||
| HT | MT | HT | MT | HT | MT | ||
| Subject 1 | ronrup | 9.66 | 9.73 | 9.60 | 9.60 | 9.73 | 9.86 |
| rupron | 9.46 | 9.66 | 9.46 | 9.53 | 9.46 | 9.33 | |
| Subject 2 | ronrup | 9.66 | 9.66 | 9.33 | 9.60 | 9.66 | 9.73 |
| rupron | 8.13 | 9.0 | 7.73 | 8.73 | 8.2 | 8.93 | |
@@
Similarly to NLLB, we have the same objective of translating between Aromanian, Romanian, and English, so we train the LLMs with samples from all six possible directions. In the case of the four models, fine-tuning for more than epoch leads to an increased loss on the dev split. Thus, we keep only the checkpoints fine-tuned for epoch with sample packing Krell et al. (2021). An example of a prompt for the LLaMA 3.1 model can be found in Table 3. More details about each model’s training hyper-parameters and prompt format are available in Appendix E.
@@ 6.1 Automatic Evaluation
We evaluate all the models in all possible directions between Aromanian, Romanian, and English. We note here that the English references are machine-translated and the results involving English should be taken with reservation. However, we provide the evaluation scores here for completeness. The automatic metrics that we employ are BLEU Post (2018), ChrF++ Popović (2015), and BERTScore Zhang et al. (2020). We report the ChrF++ scores for all the models in Table 4. All other results are available in Appendix F and Appendix G
Overall, for English-Romanian pairs, GPT-4o obtains the strongest scores in both directions, even when translating from noisy synthetic English into Romanian. Chances that the model has been previously exposed to the out-of-domain test set are very small, since the Writings are not available online. Given the synthetic nature of the English data, we cannot draw any conclusions with respect to translation quality into English.
More importantly, the scores for the Aromanian-Romanian language pairs obtained by any of the trained models are considerably higher than those obtained by GPT-4o. Among the trained models, the NLLB family consistently obtains higher scores than their fine-tuned instruction LLM counterparts, regardless of the automatic metric (see Table 4 and Appendix F).
When translating into Aromanian (rup), the differences are noticeable – in every single case NLLB-type models perform better than LLM-based translation models. Still, when translating into a well-resourced language such as English or Romanian, the performance gap narrows, with the models differing by only a few points. In-domain evaluation on the dev set (in Appendix G) shows significantly higher scores than out-of-domain evaluation on test set, with points in ChrF and double the BLEU scores. On the dev set, the differences between NLLB models and LLMs are less pronounced.
The evaluation scores do not indicate that Aromanian translation benefits more from using RoLLaMA, a model trained on Romanian instructions, or TowerInstruct v0.2, a model designed specifically for MT tasks.
All types of models struggle to generate high-quality Aromanian output, as evidenced by the overall scores. The differences range from 5 to 15 points, with translations into Aromanian scoring lower than translations from Aromanian. Appendix F provides additional metrics that align with our findings using ChrF++. We note that Costa-jussà et al. (2022) found that a increase in ChrF++ is almost always noticeable by human evaluators.
Manual inspection of translations reveals that the models tend to translate sentences word-for-word, which misses word overlaps in BLEU-based metrics for out-of-domain evaluation on mixed-genre literary texts, news, and essays. More details on manual evaluation are presented in the next section.
@@ 6.2 Human Evaluation
Two human evaluators participated in this study. Both are native Aromanian speakers from Romania, fluent in both Romanian and Aromanian. They self-identify as speakers of the Aromanian Grămustean and Cipan dialects, classified as Type A in the typology of Caragiu-Marioțeanu (1997) and Pascaru and Kahl (2017), which is also the main Aromanian variety in the dataset.
We do not evaluate for English here. The evaluated sentences contain human and machine translations in equal proportions, and therefore each annotator evaluates a total of samples. For MT, we use our NLLB 600M fine-tuned model.
Annotators are instructed to assign direct assessment scores Graham et al. (2013) from 1 to 10 to evaluate the quality of each translation (1 being the lowest score, 10 the highest). This scale is culturally motivated by the fact that it is familiar to the study participants because it is used in Romanian schools at all levels of post-primary school.
They assign a score taking into account three categories: fluency, style, and meaning (logical sense). Fluency refers to how grammatically correct the sentence is. Style denotes how likely the speaker is to build the sentence they are evaluating in the exact same way. Meaning alludes to the logical sense of the sentence.
The evaluation was carried out bilingually for all language pairs. The results are in Table 5, showing that machine translation is rated slightly better than human translation. For each low grade, we ask annotators to provide comments with a description of errors. Lexical errors predominate: annotators identify words that do not exist or are not common in their language variety in both human and machine translated texts, e.g., usage of bus - instead of aftuchină (En. bus); usage of sala di conțertu (En. concert hall); inițiativă (En. initiative) are not recognized as valid Aromanian.
Evaluating machine translation systems for nonstandard language varieties presents several challenges regarding representativeness and meaningfulness: 1) annotators may produce inconsistent results, as they may not be familiar with the full range of Aromanian varieties Pascaru and Kahl (2017); 2) machine translation systems generate neologisms derived from Romanian, even when established Aromanian terms exist, particularly when translating contemporary texts or news content; 3) human translations may receive lower ratings from speakers due to dialectal differences.
Therefore, the results of human evaluation for fine-grained quality assessment remain inconclusive. Nonetheless, the fact that of machine translations into and from Aromanian received a perfect score from human annotators suggests that the MT system has certain strengths.
@@ 7 Online Translation System
To make the machine translation model accessible to the public and raise awareness of the endangered status of the Aromanian language, we have deployed the system online. The user interface is designed to be lightweight and includes accessibility features such as copying text content and switching between the source and target languages (Figure 3).
| ronrup | rupron | |||
| Metric | BLEU | ChrF++ | BLEU | ChrF++ |
| FP32 | 17.3 | 45.03 | 30.95 | 51.01 |
| INT8 quantization | 17.01 | 44.84 | 30.21 | 50.16 |
@@ 7.1 Quantization
The fine-tuned NLLB-600M model demonstrates the highest performance and has the smallest number of parameters, making it suitable for deployment on a CPU. We quantize the model using the ctranslate2 Klein et al. (2020) engine to 8-bit integer weights. Furthermore, the engine applies additional optimizations, including layer fusion, padding removal, batch reordering, and more.
Regarding the performance loss typically associated with quantization, we observe minimal differences in the automatic metrics used. As shown in Table 6, which analyzes the Aromanian-Romanian direction on the test split, the differences are measured at less than BLEU point. Since the models are fine-tuned for sentence-level translations, the system splits user input into sentences and processes them in parallel. The system achieves a translation speed of approximately tokens per second.
7.2 Orthographic Converter
@@ 8 Conclusions
Building a dialectal machine translation system for a low-resource language comes with several challenges – from the lack of standardization and small online presence of the language to the diverse varieties specific to different countries.
Previous experiments show that modern neural machine translation systems are unable to learn from extremely small data sets of 3,000 pair of sentences, even for languages that are part of the same dialectal continuum such as Romanian and Aromanian Petrariu and Nisioi (2024). In our work, we provide the steps to construct a parallel corpus for Aromanian, which further requires training a multilingual sentence embedder. We show that a diverse parallel dataset of k examples is sufficient for fine-tuning pre-trained transformer models to do low-resource and dialectal machine translation. The preexisting knowledge of Romanian in both LLM transformers and sequence-to-sequence models enhances generalization and knowledge transfer capabilities. However, it also introduces bias and raises concerns about representativeness, as the generated Aromanian shows strong Romanian influence and similarities.
The introduction of English synthetic data in the training phase proves useful to extend the translation directions beyond Romanian. However, future work would require measuring the impact of translation errors in synthetic data.
Based on automatic evaluation, the smaller seq-to-seq models are better translators than LLM-based English-centric or multilingual fine-tuned models (that have not been pre-trained on Romanian). NLLB-based models outperform the LLM-based pre-trained models in terms of automatic metrics and have the advantage of being light-weight and easy to deploy on CPU using int8 quantization.
Regarding human evaluation, due to the relatively small size of the community and lack of access to a large pool of native speakers, we had difficulties conducting an extensive human evaluation. Furthermore, we identify several difficulties in conducting human evaluations on low-resource, unstandardized dialectal varieties, and future studies are mandatory to quantify the limitations of our MT system geographically and across language varieties.
With this work, we hope to bring Aromanian into the spotlight of machine translation and corpus-based linguistic research and contribute to preservation efforts for this endangered language.